OpenStudy (anonymous):

Simplify. √24x^8y^4 How do i find this? Can you give steps please?

OpenStudy (anonymous):

since \(24=4\times 6\) we know \(\sqrt{24}=\sqrt{4\times 6}=\sqrt{4}\times \sqrt{6}=2\sqrt{6}\)

OpenStudy (anonymous):

also \(\sqrt{x^8}=x^4\) since \((x^4)^2=x^8\) similarly \(\sqrt{y^4}=y^2\)

OpenStudy (anonymous):

so your final answer is \[2x^4y^2\sqrt{6}\]

OpenStudy (anonymous):

Thank you soo much!!